OpenStudy (baby456):

21. F(x)=1/2x; g(x)=1/3x find f(g(x) and g(f(x) State domain of each Precalc questions

OpenStudy (baby456):

@zepdrix

OpenStudy (baby456):

@Nnesha

OpenStudy (baby456):

@Owlcoffee

OpenStudy (baby456):

@leonardo0430

OpenStudy (3mar):

May I answer?

OpenStudy (3mar):

\[f(x)=\frac{ 1 }{ 2x }\] and \[g(x)=\frac{ 1 }{ 3x }\] To get f(g(x)), replace every x with g(x): \[f(g(x))=\frac{ 1 }{ 2(\frac{ 1 }{ 3x }) }=\frac{ 3x }{ 2 }\] Its domain is all available possible values of x, i.e all real numbers. Similarly, to get g(f(x) you replace every x with f(x): \[g(f(x))=\frac{ 1 }{ 3(\frac{ 1 }{ 2x }) }=\frac{ 2x }{ 3 }\] Its domain is all real numbers.

OpenStudy (3mar):

Thanks for the medal!